OpenStudy (ray1998):

Please help? Will give fan and medal! The weight P in pounds that a beam can safely carry is inversely proportional to the distance D in feet between the supports of the beam. For a certain type of wooden beam, P = 9,200/D. What distance between supports is needed to carry 2000 lb? For the beam to support a weight of 2,000 lb, the distance between the supports needs to be no more than ___ ft. Thank you all in advance!!

OpenStudy (babalooo):

Hmmm...do you have a answer key to this? I'm thinking P=9200/D P=2000=9200/D 2000D=9200 D=9200/2000=4.6 Isn't that it?

OpenStudy (ray1998):

That actually is it, THANK YOU SOOOO MUCH!!! I've been working on this problem for hours. (I'm very slow at math) and finally just gave up and came here. Thank you so much for the help :) I really appriciate it. Normally I'd ask more questions about it and have the person go over it with me, but I'm too sleep deprived to care lol so again, thank you very much. @babalooo , sorry for the longish reply ._.

OpenStudy (babalooo):

no worries. best of luck with your homework ;D

OpenStudy (ray1998):

Thank you ^.^